Fast Turnstiles Co. is evaluating the extension of credit to a new group of customers. Although these customers will provide $234,000 in additional credit sales, 15 percent are likely to be uncollectible. The company will also incur $16,500 in additional collection expense. Production and marketing costs represent 70 percent of sales. The firm is in a 30 percent tax bracket and has a receivables turnover of four times. No other asset buildup will be required to service the new customers. The firm has a 10 percent desired return.  

a-1. Calculate the incremental income after taxes.  
  

    

a-2. Calculate the return on incremental investment. (Input your answer as a percent rounded to 2 decimal places.)
  

   

a-3. Should Fast Turnstiles Co. extend credit to these customers?
  

Yes
No

  

b-1. Calculate the incremental income after taxes if 18 percent of the new sales prove to be uncollectible.
  

   

b-2. Calculate the return on incremental investment if 18 percent of the new sales prove to be uncollectible. (Input your answer as a percent rounded to 2 decimal places.)
  

     

b-3. Should credit be extended if 18 percent of the new sales prove uncollectible?
  

Yes
No

  

c-1. Calculate the return on incremental investment if the receivables turnover drops to 1.6, and 15 percent of the accounts are uncollectible. (Input your answer as a percent rounded to 2 decimal places.)
  

   

c-2. Should credit be extended if the receivables turnover drops to 1.6, and 15 percent of the accounts are uncollectible?
Answer #1

a-1. Incremental income after taxes : $ 13,020.

Increase in Sales $ 234,000
Less: Accounts Uncollectible 35,100
Annual Incremental Revenue 198,900
Less: Collection Expense 16,500
Less: Production and Marketing Costs 163,800
Income before taxes 18,600
Tax Expense ( 30 % ) 5,580
Net Income 13,020

a-2. Return on Incremental Investment = 22.26 %

Investment in Accounts receivable = Incremental Sales / Receivables Turnover = $ 234,000 / 4 = $ 58,500.

Return on incremental investment = $ 13,020 / $ 58,500 * 100 = 22.26 %

a-3. Yes.

Since the return on incremental investment of 22.26 % exceeds the firm's desired return of 10 %.

b-1. Incremental income after taxes : $ 8,106

Increase in Sales $ 234,000
Less:Accounts Uncollectible 42,120
Incremental Sales Revenue $ 191,880
Less: Collection Expense 16,500
Less: Production and Marketing Expense 163,800
Income before taxes 11,580
Income Tax Expense 3,474
Incremental Income after Taxes 8,106

b-2: Return on incremental investment : 13.86 %

b-3. Yes.

c-1.Return on Incremental Investment : 8.90 %

Investment in accounts receivable = $ 234,000 / 1.6 = $ 146,250.

Return on incremental investment = $ 13,020 / $ 146,250 * 100 = 8.90 %

c-2. No.